# Diffusion approximation of the stationary distribution of a two-level single server queue

### v7

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study the limit of the stationary distribution of the queue length in this two-level queue under scaling in heavy traffic. This limit corresponds to its diffusion approximation. It is shown that this limiting distribution is truncated exponential (or uniform if the drift is null) below the threshold level and exponential above it under suitably chosen system parameters and generally distributed inter-arrival times and workloads brought by customers. This result is proved under a mild limitation on arrival parameters using the so called BAR approach studied in [3, 4, 10, 12]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distribution under scaling.
@@ 1 Introduction
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length. Namely, customers arrive subject to a renewal process as long as its queue length (including a customer being served) is less than threshold $\ell_{0}$, while it is changed to another renewal process otherwise. The service rate is changed similarly but in a slightly different way. That is, the service rate is constant when the queue length is not greater than the threshold, while it is another constant otherwise. We assume that customers are served in the FCFS manner, and the amounts of the work which are brought by them are independent and identically distributed and independent of the arrival process. We refer to this model as a 2-level $GI/G/1$ queue.
@@
This 2-level $GI/G/1$ queue is motivated by an energy saving problem to control the arrival process and service rate. Here, the system is an internet server, and its energy consumption depends on service rate. Then, the problem is to find an optimal threshold to minimize the operating energy of the server and customer’s waiting cost.
@@
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available (see Appendix A.2), but the variability of the arrival times and service amounts are not well captured in this exponential case. Because of these facts, an approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution. This limiting distribution corresponds to the stationary distribution of a one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold, which is recently studied in [11, 13]. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
@@
There is another motivation of this study. In general, theoretical study of diffusion approximation is hard for state-dependent queues. For example, there are some theoretical studies about them, but they assume that the arrival and service processes are of exponential types, and their rates are smoothly changed (e.g., see [8, 15]), while the arrival process and service rate discontinuously change in the present model. Thus, theoretical study for its diffusion approximation is challenging.
@@
As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue in heavy traffic (e.g., see [7]). Furthermore, [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room. Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see [6]). Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it. This conjecture can be directly verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional. Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [3, 4, 10, 12]. This verification, Theorem 3.1, requires a mild limitation on the parameters of the arrival processes (see (3.f)). Its proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory. However, the proof has a complicated structure, so we split it into many components so that they are orthogonal as long as possible. This is the reason why we spend two sections for the proof.
@@
In this paper, we only study the limit of the scaled stationary distribution, but, as our results are related to the diffusion approximation, it is interesting to identify a reflected Brownian motion as a process limit for the 2-level $GI/G/1$ queue. This is a hard question because this Brownian motion discontinuously changes its drift and variance at the fixed state. We intuitively discuss what can be expected about this process limit in Section 6.1.
@@ 2 Framework for the 2-level $GI/G/1$ queue
We take the stochastic framework introduced in [12] to study the 2-level $GI/G/1$ queue. Let $L(t)$, $R_{e}(t)$ and $R_{s}(t)$ be the queue length, the remaining arrive time and the remaining workload of a customer being served, respectively, at time $t$. Note that the remaining service time cannot be used for $R_{s}(t)$ here because the service rate may change in future. Define
@@
According to [12], we take stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}\equiv\{\Theta_{t};t\geq 0\}$ to jointly consider the Markov process $X(\cdot)$ and the counting processes $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where $\mathbb{F}\equiv\{{\mathcal{F}}_{t};t\geq 0\}$ is a filtration generated by $X(\cdot)$, that is, ${\mathcal{F}}_{t}=\sigma(\{X(u);0\leq u\leq t\})$, and $\{\Theta_{t};t\geq 0\}$ is a operator semi-group (namely, $\Theta_{s+t}=\Theta_{s}\circ\Theta_{t}$ for $s,t\geq 0$) on $\Omega$ satisfying the following two conditions.
@@ 3 A sequence of the 2-level $GI/G/1$ queues
We consider the scaling limit of the stationary distribution of the 2-level $GI/G/1$ queue in heavy traffic. To this end, we introduce a sequence of this queueing models indexed by $r\in(0,1]$. Since this sequence can be countable, one may take $r=1/\sqrt{n}$ for positive integer $n$ so that the scaling is identical with the standard diffusion approximation, but we keep $r$ just for convenience.
@@ 3.1 Assumptions
For the $r$-th model, let $\ell^{(r)}_{0}\geq 2$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$. We refer to this number as the queue length. Let $t^{(r)}_{e,n}$ be the $n$-th arrival time of a customer, and let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customer. We assume that $T_{e}^{(r)}(n)\equiv t^{(r)}_{e,n}-t^{(r)}_{e,n-1}$ is independent of $\{t^{(r)}_{e,n^{\prime}};0\leq n^{\prime}\leq n-1\}$ and everything else for all $n\geq 1$, where $t^{(r)}_{e,0}=0$. Thus, the inter-arrival times are independent, but they may have two different distributions depending on the queue length. Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are independent random variables which may have different distributions. As for service, let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customers. We assume that $\{T^{(r)}_{s}(n);n\geq 1\}$ is a sequence of $i.i.d.$ random variable. We denote a random variable subject to the common distribution of $T^{(r)}_{s}(n)$ by $T^{(r)}_{s}$. Let $c^{(r)}_{1}$ and $c^{(r)}_{2}$ be the service speeds when the queue length is not greater than $\ell^{(r)}_{0}$ and otherwise, respectively. Namely, as long as $R^{(r)}_{s}(t)$ is differentiable at time $t$,
$\frac{d}{dt}R^{(r)}_{s}(t)=\left\{\begin{array}[]{ll}-c^{(r)}_{1},&L^{(r)}(t)\leq\ell^{(r)}_{0},\\ -c^{(r)}_{2},&L^{(r)}(t)>\ell^{(r)}_{0}.\end{array}\right.$
These dynamics is depicted in the sample path of $L^{(r)}(t)$ below.
Figure 3.1: Sample path of the queue length process
@@
$m^{(r)}_{e,i}=\mathbb{E}(T^{(r)}_{e,i}),\qquad\lambda^{(r)}_{i}=1/m^{(r)}_{e,i},\qquad\sigma^{(r)}_{e,i}=\mathbb{E}[(T^{(r)}_{e,i}-m^{(r)}_{e,i})^{2}],\qquad i=1,2,$
$m^{(r)}_{s}=\mathbb{E}(T^{(r)}_{s}),\qquad\mu^{(r)}=1/m^{(r)}_{s},\qquad\sigma^{(r)}_{s}=\mathbb{E}[(T^{(r)}_{s}-m^{(r)}_{s})^{2}],$
@@
For the $r$-th model, at time $t$, let $R^{(r)}_{e}(t)$ be the remaining arrival time, and let $R^{(r)}_{s}(t)$ be the remaining workload of a customer being served. Define $X^{(r)}(\cdot)\equiv\{X^{(r)}(t);t\geq 0\}$ by
@@
and denote the counting processes of the arrivals and service completions by $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$, respectively. Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov. Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with the time-shift operator $\Psi$. We assume the following assumptions.
@@
    For $i=1,2$, $m^{(r)}_{e,i},\lambda^{(r)}_{i},\sigma^{(r)}_{e,i},c^{(r)}_{i}$ converge to $m_{e,i},\lambda_{i},\sigma_{e,i},c_{i}>0$, respectively, as $r\downarrow 0$. Similarly, $m^{(r)}_{s},\mu^{(r)},\sigma^{(r)}_{s}$ converge to $m_{s},\mu,\sigma_{s}$, respectively, as $r\downarrow 0$.
@@
    There is a constant $\ell_{0}$ such that $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$.
@@
Here, for positive valued functions $f,g$ of $r$, we write $f(r)=o(g(r))$ as $r\downarrow 0$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$. Similarly, $f(r)=O(g(r))$ if $\limsup_{r\downarrow 0}f(r)/g(r)<\infty$. Note that (3.a) implies that
@@
The condition (3.f) comes out from technical requirements for the proof of Theorem 3.1. It is a mild limitation because no condition is needed for $\lambda_{1}\not=\lambda_{2}$, but we think that it is likely not needed. However, removing (3.f) seems to be beyond the framework of this paper. See Remark 5.1 for its details.
We take countably many $r$’s for a limiting operation. For such a sequence, there exists a stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator $\Theta_{\bullet}$ on which $X^{(r)}(\cdot)$ is a stationary process. In what follows, $\mathbb{P}$ is always this probability measure. By the condition (3.e), the Markov process $X^{(r)}(\cdot)$ has the stationary distribution. Let $X^{(r)}\equiv(L^{(r)},R^{(r)}_{e},R^{(r)}_{s})$ be a random vector subject to the stationary distribution of $X^{(r)}(t)$. For convenience, let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$. For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process obtained from the renewal process whose inter-counting times have the same distribution as that of $T^{(r)}_{e,i}$.
Define $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]$ and $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]$, and let $a\wedge b=\min(a,b)$ for $a,b\in\mathbb{R}$, which will be used throughout the paper. We note the following simple but important facts.
@@
Assume the assumptions (3.a)–(3.e) and the stationarity of $X^{(r)}(\cdot)$, then $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite.
Proof.
Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$. Obviously, $\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,0}(t)]$ under the stationary framework. Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}\wedge T_{e,2}]<\infty$. Similarly, let $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $((c^{(r)}_{1})^{-1}\wedge(c^{(r)}_{1}))^{-1}T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[((c^{(r)}_{1})^{-1}\wedge(c^{(r)}_{1})^{-1})T_{s}]<\infty$. ∎

### v8

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service speed by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service speed are controlled. To get its performance in tractable form, we study the limit of the stationary distribution of the queue length in this two-level queue under scaling in heavy traffic. Except for a special case, this limit corresponds to its diffusion approximation. It is shown that this limiting distribution is truncated exponential (or uniform if the drift is null) below the threshold level and exponential above it under suitably chosen system parameters and generally distributed inter-arrival times and workloads brought by customers. This result is proved under a mild limitation on arrival parameters using the so called BAR approach studied in [3, 4, 10, 12]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distribution under scaling.
@@ 1 Introduction
We consider a single server queue which has a threshold to change its arrival process and g by its queue length. Namely, customers arrive subject to a renewal process as long as its queue length (including a customer being served) is less than threshold $\ell_{1}$, while it is changed to another renewal process otherwise. The service speed is changed similarly but in a slightly different way. That is, the service speed is constant when the queue length is not greater than the threshold, while it is another constant otherwise. We assume that customers are served in the FCFS manner, and the amounts of the work which are brought by them are independent and identically distributed and independent of the arrival process. We refer to this model as a 2-level $GI/G/1$ queue.
@@
This 2-level $GI/G/1$ queue is motivated by an energy saving problem to control the arrival process and service speed. Here, the system is an internet server, and its energy consumption depends on service speed. Then, the problem is to find an optimal threshold to minimize the total cost of the operating energy of the server and customer’s waiting.
@@
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available, but the variability of the arrival times and service amounts are not well captured in this exponential case. Because of these facts, an approximation by a reflected diffusion is helpful, where a reflected diffusion means a one-dimensional diffusion reflected at the origin, in which the term “diffusion” is used for a one-dimensional continuous Markov process on the real line which has a finite and positive second moment at each time.
So we study the diffusion approximation through the limit of the stationary distributions of the scaled queue length of this 2-level $GI/G/1$ queue in heavy traffic, where the scaling factor and threshold level are chosen so that the limit is well obtained as a probability distribution. Except for a special case, this limiting distribution corresponds to the stationary distribution of a reflected diffusion whose drift and variance are discontinuously changed at the level corresponding to the threshold (see Section 6.1). The stationary distribution of this reflected diffusion is recently studied in [11, 13]. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
@@
There is another motivation of this study. In general, it is hard to theoretically study diffusion approximation for state-dependent queues. For example, there are some theoretical studies about them, but they assume that the arrival and service processes are of exponential types, and their rates are smoothly changed (e.g., see [8, 15]), while the arrival process and service speed discontinuously change in the present model. Thus, theoretical study for its diffusion approximation is challenging.
@@
As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue in heavy traffic (e.g., see [7]). Furthermore, [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room. Both distributions agree with the stationary distributions of the corresponding reflected diffusions (see [6]). Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential or uniform below the threshold level and exponential above it. This conjecture can be directly verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional. Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [3, 4, 10, 12]. This verification, Theorem 3.1, requires a mild limitation on the parameters of the arrival process (see (3.f) in Section 3.1). Its proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory. However, the proof is complicated, so we split it into several components so that they are orthogonal as long as possible. This is the reason why we spend two sections for the proof.
@@
In this paper, we only study the limit of the scaled stationary distribution, but, as our results are related to the diffusion approximation, it is interesting to identify a reflected diffusion as a process limit for the 2-level $GI/G/1$ queue. This is a hard question because this diffusion discontinuously changes its drift and variance at the fixed state. We intuitively discuss what can be expected about this process limit in Section 6.1.
@@ 2 Framework for the 2-level $GI/G/1$ queue
We take the stochastic framework introduced in [12] to study the 2-level $GI/G/1$ queue. Let $L(t)$, $R_{e}(t)$ and $R_{s}(t)$ be the queue length, the remaining arrive time and the remaining workload of a customer being served, respectively, at time $t$. Note that the remaining service time cannot be used for $R_{s}(t)$ here because the service speed may change in future. Define
@@
According to [12], we take stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}\equiv\{\Theta_{t};t\geq 0\}$ to jointly consider the Markov process $X(\cdot)$ and the counting processes $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where $\mathbb{F}\equiv\{{\mathcal{F}}_{t};t\geq 0\}$ is a filtration generated by $X(\cdot)$, that is, ${\mathcal{F}}_{t}=\sigma(\{X(u);0\leq u\leq t\})$, and $\Theta_{\bullet}$ is a operator semi-group (namely, $\Theta_{s+t}=\Theta_{s}\circ\Theta_{t}$ for $s,t\geq 0$) on $\Omega$ satisfying the following two conditions.
@@ 3 A sequence of the 2-level $GI/G/1$ queues
We consider the scaling limit of the stationary distribution of the 2-level $GI/G/1$ queue in heavy traffic. To this end, we introduce a sequence of this queueing models indexed by $r\in(0,1]$. Since this sequence can be countable, one may take $r=1/\sqrt{n}$ for positive integer $n$ so that the scaling is identical with the standard diffusion approximation, but we keep $r$ just for convenience. Let $a\wedge b=\min(a,b)$ and $a\vee b=\max(a,b)$ for $a,b\in\mathbb{R}$, which will be used throughout the paper.
@@ 3.1 Assumptions
For the $r$-th model, let $\ell^{(r)}_{1}\geq 1$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$. We refer to this number as the queue length. Let $t^{(r)}_{e,n}$ be the $n$-th arrival time of a customer, and let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customer. We assume that $T_{e}^{(r)}(n)\equiv t^{(r)}_{e,n}-t^{(r)}_{e,n-1}$ is independent of $\{t^{(r)}_{e,n^{\prime}};0\leq n^{\prime}\leq n-1\}$ and everything else for all $n\geq 1$, where $t^{(r)}_{e,0}=0$. Thus, the inter-arrival times are independent, but their distributions depend on the queue length. Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{1}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{1}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are independent random variables which may have different distributions. As for service, let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customers. We assume that $\{T^{(r)}_{s}(n);n\geq 1\}$ is a sequence of $i.i.d.$ random variables. We denote a random variable subject to the common distribution of $T^{(r)}_{s}(n)$ by $T^{(r)}_{s}$. Let $c^{(r)}_{1}$ and $c^{(r)}_{2}$ be the service speeds when the queue length is not greater than $\ell^{(r)}_{1}$ and otherwise, respectively. Namely, as long as $R^{(r)}_{s}(t)$ is differentiable at time $t$,
$\frac{d}{dt}R^{(r)}_{s}(t)=-c^{(r)}_{1}1(1\leq L^{(r)}(t)\leq\ell^{(r)}_{1})-c^{(r)}_{2}1(L^{(r)}(t)>\ell^{(r)}_{1}),$
where $1(\cdot)$ is the indicator function of the statement $\cdot$”. These dynamics is depicted in the sample path of $L^{(r)}(t)$ below.
Figure 3.1: Sample path of the queue length process of the $r$-th model
@@
$m^{(r)}_{e,i}=\mathbb{E}(T^{(r)}_{e,i}),\qquad\lambda^{(r)}_{i}=1/m^{(r)}_{e,i},\qquad\sigma^{(r)}_{e,i}=(\mathbb{E}[(T^{(r)}_{e,i}-m^{(r)}_{e,i})^{2}]^{-\frac{1}{2}},\qquad i=1,2,$
$m^{(r)}_{s}=\mathbb{E}(T^{(r)}_{s}),\qquad\mu^{(r)}=1/m^{(r)}_{s},\qquad\sigma^{(r)}_{s}=(\mathbb{E}[(T^{(r)}_{s}-m^{(r)}_{s})^{2}])^{-\frac{1}{2}},$
@@
For the $r$-th model, let $R^{(r)}_{e}(t)$ and $R^{(r)}_{s}(t)$ be the remaining arrival time and the remaining workload of a customer being served, respectively, at time $t$. Define $X^{(r)}(\cdot)\equiv\{X^{(r)}(t);t\geq 0\}$ by
@@
and denote the counting processes of the arrivals and service completions by $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$, respectively. Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov (e.g., see [5]). Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with the time-shift operator $\Theta_{\bullet}$. We assume the following assumptions.
@@
    For $i=1,2$, $m^{(r)}_{e,i},\lambda^{(r)}_{i},\sigma^{(r)}_{e,i},c^{(r)}_{i}$ and $m^{(r)}_{s},\mu^{(r)},\sigma^{(r)}_{s}$ are finite and positive, and converge to $m_{e,i},\lambda_{i},\sigma_{e,i},c_{i}>0$ and $m_{s},\mu,\sigma_{s}>0$, respectively, as $r\downarrow 0$.
@@
    There is a constant $\ell_{1}$ such that $r\ell^{(r)}_{1}=\ell_{1}+o(r)$ as $r\downarrow 0$.
@@
Here, for functions $f,g:(0,1]\to(0,\infty)$, we write $f(r)=o(g(r))$ as $r\downarrow 0$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$. Similarly, $f(r)=O(g(r))$ if $\limsup_{r\downarrow 0}f(r)/g(r)<\infty$. Note that (3.a) implies that
@@
This is a mild limitation because no condition is needed for $\lambda_{1}\not=\lambda_{2}$ and it is automatically satisfied when $T_{e,1}$ and $T_{e,2}$ are identically distributed. This condition is likely not needed but hard to remove. See Remark 5.1 for its details.
We take countably many $r$’s for a limiting operation. For such a sequence, there exists a stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator $\Theta_{\bullet}$ on which $X^{(r)}(\cdot)$ is a stationary process. In what follows, $\mathbb{P}$ is always this probability measure. By the condition (3.e), the Markov process $X^{(r)}(\cdot)$ has the stationary distribution. Let $X^{(r)}\equiv(L^{(r)},R^{(r)}_{e},R^{(r)}_{s})$ be a random vector subject to the stationary distribution of $X^{(r)}(t)$. For convenience, let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$. For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process obtained from the renewal process whose inter-counting times have the same distribution as that of $T^{(r)}_{e,i}$. Similarly, let $N^{(r)}_{s}$ be the stationary point process for the service completion instants.
Define $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]$ and $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]$. At present, we cannot compute $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$, but the following lemma is intuitively clear because $X^{(r)}(\cdot)$ is a stationary process by (3.e). Since this lemma is a key for our analysis, we formally prove it in Appendix A.1.
@@
Assume the assumptions (3.b) and (3.e), then $\alpha^{(r)}_{e}$ is finite and positive, and $\alpha^{(r)}_{s}=\alpha^{(r)}_{e}$.